the coefficient of static friction between block A of mass 2 kg and the table is 0.1. what would be the maximum mass of block B so that two blocks do not move? the string and the pulley are assumed to be smooth and massless? tak g =10 m per second square

Grade:10

2 Answers

Aman Bansal
12 years ago

Dear Arup,

The details given in the question are incomplete , Its not given if the other block is inclined or hanging or just placed.

Best Of luck

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Mujahid Ahmed
12 years ago

Apply Condition of no slipping 

T < 0.1* N

  Tmax   = 0.1 * 20

             = 2

 

Tmax = Mmax * g

2 = M*10

M = 0.2